Q waves on the electrocardiogram, according to Schinkel et al. (A-15), are often considered to be

reflective of irreversibly scarred myocardium. These researchers assert, however, that there are some indications that residual viable tissue may be present in Q-wave-infarcted regions. Their study of 150 patients with chronic electrocardiographic Q-wave infarction found 202 dysfunctional Q-wave regions. With dobutamine stress echocardiography (DSE), they noted that 118 of these 202 regions were viable with information from the DSE testing. Construct a 90 percent confidence interval for the proportion of viable regions that one might expect to find a population of dysfunctional Q-wave regions.



1
Expert's answer
2021-07-16T09:07:34-0400

The proportion of region expected

"p = \frac{118}{202} = 0.584 \\

n = 202"

The z score for 90% confidence level = 1.645

Confidence interval formula:

"p±Z \times \sqrt{ \frac{p(1-p)}{n} }"

The margin of error "= 1.645 \times \sqrt{ \frac{0.584 \times 0.416}{202} }=0.057"

The confidence interval = 0.584 ± 0.057

= (0.527, 0.641)